OpenStudy (anonymous):

(g + 4) – 3g = 1 + g

Parth (parthkohli):

Remove the brackets. People suck when they put it and there's a positive sign outside :P

OpenStudy (jiteshmeghwal9):

\[g+4=1+g+3g\] Now u can solve it, can't u?

OpenStudy (jiteshmeghwal9):

\[g+4=1+4g\]

OpenStudy (anonymous):

take all the g in right hand side and no. on left or reverse

OpenStudy (jiteshmeghwal9):

\[4-1=4g-g\]

OpenStudy (jiteshmeghwal9):

\[3=3g\]

OpenStudy (anonymous):

g=1

OpenStudy (jiteshmeghwal9):

Therefore,\[g=1\]

OpenStudy (jiteshmeghwal9):

i think now u can get it:)

OpenStudy (anonymous):

i have a advice make the question as easy you can first open the brackets and all that

OpenStudy (anonymous):

for ex. a equation \[(2x ^{7})^{2}+91\]

OpenStudy (anonymous):

it is hard then simply it \[2x ^{14}+91\]

OpenStudy (anonymous):

like this